A small fly of mass 0.22 g is caught in a spider’s web. The web oscillates predominantly with a frequency of 4.0 Hz. (a) What is the value of the effective spring stiffness constant k for the web? (b) At what frequency would you expect the web to oscillate if an insect of mass 0.44 g were trapped?


1
Expert's answer
2021-03-14T19:12:50-0400

(a) We can find the value of the effective spring stiffness constant "k" for the web as follows:


"f=\dfrac{1}{2\pi}\sqrt{\dfrac{k}{m}},""k=4\pi^2f^2m,""k=4\pi^2\cdot(4.0\ Hz)^2\cdot0.22\cdot10^{-3}\ kg=0.14\ \dfrac{N}{m}."

(b) We can find the frequency of oscillations of the web as follows:


"f=\dfrac{1}{2\pi}\sqrt{\dfrac{0.14\ \dfrac{N}{m}}{0.44\cdot10^{-3}\ kg}}=2.84\ Hz."